OpenStudy (anonymous):

what is the diameter of the circle if its radius is: 1) 10 cm 2) 6.7 m 3) 25 dm 4) 9.5 ft 5) (4x+3) cm

OpenStudy (anonymous):

pls answer this pls.......

OpenStudy (anonymous):

The diameter is just the radius times two. 1) 20 cm 2) 13.4 m 3) 50 dm 4) 19 ft 5) (8x+6) cm

OpenStudy (anonymous):

thank you.... yellowranger 15

OpenStudy (anonymous):

how about the 3\[\sqrt{7}\] m

OpenStudy (anonymous):

\[3\sqrt{7} m\]

OpenStudy (anonymous):

That would be \[6\sqrt{7}m\]